# Variational Transformer Ansatz for the Density Operator of Steady States in Dissipative Quantum Many-Body Systems

### v1

@@ Abstract
The experimental results, presented in Fig. 5, compare the performance of our proposed transformer density operator ansatz with the exact values. The exact curve is calculated by exact diagonalization as described in III.6 in the supplementary material. As shown in the graph, the result from our ansatz matches the exact curve well. These results highlight the robustness of the transformer density operator ansatz in capturing the complex features of the spin system across varying magnetic field strengths.
Conclusion and discussion. — In this work, we introduce a transformer variational ansatz to efficiently encode and solve the steady states of dissipative quantum systems. By vectorizing the density operator into a many-body state, we first embed the input configurations into feature vectors and then utilize the self-attention mechanisms to model long-range correlation, which is crucial in the open system. Numerical experiments on paradigmatic models, such as the dissipative transverse-field Ising model and the Heisenberg model, show that our approach accurately reproduces steady state of various dissipative systems while maintaining a compact parameterization and achieving fast convergence.
@@
Acknowledgments. — We acknowledge Di Luo, Filippo Vicentini, Yuan-Hang Zhang, and Chen Zhuo for beneficial communications. We thank Filippo Vicentini and Di Luo for sharing their codes with us. The numerical implementation of the variational transformer density operator ansatz was done using JAX. The variational quantum Monte Carlo and stochastic reconfiguration optimizers are available in NetKet. Z. J. and D. K. are supported by the National Research Foundation in Singapore and A*STAR under its CQT Bridging Grant and CQT- Return of PIs EOM YR1- 10 Funding.
Supplementary Material: Variational Transformer Ansatz for the Density Operator of Steady States in Dissipative Quantum Many-Body Systems
In this supplementary material, we provide a detailed discussion of our transformer density operator ansatz for solving the steady state. In Section I, we discuss the vectorized Lindblad equation, and in Section II and Section III, we provide an in-depth description of our transformer density operator ansatz and the optimization machanism.
I Steady state and the vectorization of density operator
In this section, we review the vectorization formalism of the density operator and its dynamics, a powerful framework for representing the steady states of dissipative open quantum systems. This approach has recently gained significant attention in studies of open-system quantum phases, weak and strong symmetries, tenfold classification, and other related topics (see, e.g., Kawabata et al. (2023); Bao et al. (2023); Sohal and Prem (2025); Yoshioka and Hamazaki (2019); Ma and Turzillo (2024)). Unlike the traditional density matrix representation, the vectorized form offers a more convenient computational framework for steady-state analysis. By expressing the density operator in this form, steady states can be obtained by solving for the ground state of a specially constructed operator.
@@ I Steady state and the vectorization of density operator
$\hat{\rho}=\sum_{\alpha,\beta}\rho_{\alpha,\beta}|\alpha\rangle\langle\beta|,$ (S1)
@@
$|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}=\sum_{\alpha,\beta}\rho_{\alpha,\beta}|\alpha\rangle|\beta\rangle.$ (S2)
@@
$|(|\psi\rangle\langle\phi|)\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}=|\psi\rangle|\phi^{*}\rangle,$ (S3)
@@
$|A\otimes B\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}\neq|A\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}\otimes|B\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}.$⟩ . (S4)
@@
$A\rho B\mapsto(A\otimes B^{T})|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}.$⟩ . (S5)
@@
$|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}=\sum_{\alpha_{1},\dots,\alpha_{N};\beta_{1},\dots,\beta_{N}}\rho_{\alpha_{1},\dots,\alpha_{N};\beta_{1},\dots,\beta_{N}}|\alpha_{1},\dots,\alpha_{N}\rangle\otimes|\beta_{1},\dots,\beta_{N}\rangle.$ (S6)
@@
$\frac{d\rho}{dt}=\mathcal{L}(\rho)={-i}[H,\rho]+\sum_{i>0}\gamma_{i}(L_{i}\rho L_{i}^{\dagger}-\frac{1}{2}\{L_{i}^{\dagger}L_{i},\rho\}),$ (S7)
@@
$\frac{d}{dt}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}=\hat{\mathcal{L}}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}},$⟩ , (S8)
@@
$\hat{\mathcal{L}}=$ $-i(H\otimes\mathds{I}-\mathds{I}\otimes H^{T})+\sum_{i>0}\gamma_{i}[L_{i}\otimes L^{*}_{i}-\frac{1}{2}(L^{\dagger}_{i}L_{i}\otimes\mathds{I}+\mathds{I}\otimes L_{i}^{T}L_{i}^{*})].$ (S9)
@@
$\mathfrak{L}=\hat{\mathcal{L}}^{\dagger}\hat{\mathcal{L}},$ (S11)
@@
provides the steady state, as shown in equation (S10). We can then optimize the energy functional to find the ground state, akin to the approach in closed systems. The energy functional is given by
$E=\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\rho|\mathfrak{L}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}=\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\rho|\hat{\mathcal{L}}^{\dagger}\hat{\mathcal{L}}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}.$⟩ . (S13)
@@
$\operatorname{Tr}(A^{\dagger}B)=\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}A|B\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}.$⟩ . (S14)
@@
$\langle O\rangle=\operatorname{Tr}(O\rho)=\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}O|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}},$⟩ , (S15)
@@
II Transformer Density Operator Architecture
@@
II.0.1 Overview of the Transformer Density Operator Architecture
@@ II.0.2 Feature Embedding
$\mathbf{X}^{(1)}\;=\;\mathrm{Conv}_{\text{circular}}\!\bigl{(}\mathbf{X}_{\mathrm{input}}\bigr{)}\;\in\;\mathbb{R}^{B\times L\times 1\times C_{1}},$ (S16)
@@
$\mathbf{X}^{(2)}\;=\;\mathrm{Conv}_{\text{circular}}\!\bigl{(}\mathbf{X}^{(1)}\bigr{)}\;\in\;\mathbb{R}^{B\times L\times 1\times C_{2}}.$ (S17)
@@
II.0.3 Self-Attention for Global Correlations
@@
$\mathbf{X}\;=\;\mathrm{squeeze}\!\bigl{(}\mathbf{X}^{(2)},\,\texttt{axis}=2\bigr{)}\;\in\;\mathbb{R}^{B\times L\times C_{2}}.$ (S18)
@@
$\mathbf{Q}$ $=\mathbf{X}\,\mathbf{W}^{Q},\quad\mathbf{K}\;=\;\mathbf{X}\,\mathbf{W}^{K},\quad\mathbf{V}\;=\;\mathbf{X}\,\mathbf{W}^{V},$ (S19)
@@
$\mathrm{head}_{i}\;=\;\mathrm{softmax}\!\Bigl{(}\frac{\mathbf{Q}_{i}\,\mathbf{K}_{i}^{\top}}{\sqrt{C_{2}/h}}\Bigr{)}\,\mathbf{V}_{i}$ (S20)
@@
$\mathbf{X}^{\prime}\;=\;\mathrm{Concat}(\mathrm{head}_{1},\dots,\mathrm{head}_{h})\,\mathbf{W}^{O}\;+\;\mathbf{X}.$ (S21)
@@
II.0.4 Global Mean-Pooling and Final Output
@@
$\mathbf{X}^{(\mathrm{pool})}\;=\;\frac{1}{L}\sum_{j=1}^{L}\bigl{(}\mathbf{X}^{\prime}_{:\,,j\,,:\,}\bigr{)},\quad\mathbf{X}^{(\mathrm{pool})}\,\in\,\mathbb{R}^{B\times C_{2}}.$ (S22)
@@
III Optimization and Evaluation of Transformer Density Operator Ansatz
@@
III.1 Efficient Sampling Strategy
@@
$|\rho_{\bm{\theta}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}=\sum_{\bm{\alpha},\bm{\beta}}\rho_{\bm{\theta}}(\bm{\alpha},\bm{\beta})|\bm{\alpha}\rangle\otimes|\bm{\beta}\rangle.$ (S23)
@@
$P_{\bm{\theta}}(\bm{\alpha},\bm{\beta})\propto|\rho_{\bm{\theta}}(\bm{\alpha},\bm{\beta})|^{2}$ (S24)
@@
    $\mathbf{A}_{\textbf{accept}}((\bm{\alpha},\bm{\beta})\to(\bm{\alpha}^{\prime},\bm{\beta}^{\prime}))=\min\left(1,~{}\frac{P_{\bm{\theta}}(\bm{\alpha}^{\prime},\bm{\beta}^{\prime})g((\bm{\alpha},\bm{\beta})|(\bm{\alpha}^{\prime},\bm{\beta}^{\prime}))}{P_{\bm{\theta}}(\bm{\alpha},\bm{\beta})g((\bm{\alpha}^{\prime},\bm{\beta}^{\prime})|(\bm{\alpha},\bm{\beta}))}\right),$ (S25)
@@
III.2 Mixed-State Observables

### v2

@@ Abstract
The experimental results, presented in Fig. 6, compare the performance of our proposed transformer density operator ansatz with the exact baseline. The exact curve (black line) is obtained by the iterative BiCGStab solver of NetKet as described in C.6 in the supplementary material. This method can give results with high accuracy for system of size larger than 7 and we utilize it as our baseline. As shown in the graph, the result from our ansatz matches the exact curve well. In the supplementary material D, we also provided the result of 10-site one-dimensional Heisenberg model. These results highlight the robustness of the transformer density operator ansatz in capturing the complex features of the spin system across varying magnetic field strengths.
Numerical results for 2D dissipative Ising model. — To demonstrate the scalability of our approach to higher-dimensional systems, we extend our study to a 2D dissipative Ising model defined on an $M\times N$ square lattice. The Hamiltonian takes the form
$H=\frac{V}{4}\sum_{\langle ij\rangle}\sigma^{z}_{i}\sigma^{z}_{j}+\frac{g}{2}\sum_{i\in\mathbb{Z}_{M}\times\mathbb{Z}_{N}}\sigma^{x}_{i},$ (17)
where the notation $\langle ij\rangle$ indicates a summation over all nearest-neighbor pairs. The local jump operator is taken to be $L_{i}=\sigma^{-}_{i}$, as discussed in, e.g., Ref. Roberts and Clerk (2023).
We investigate a $2\times 3$ lattice system with periodic boundary conditions. The system parameters are set to $V=2\gamma$ and we vary the transverse field strength $g/\gamma$ to explore different regimes of the phase diagram. For this system size, exact diagonalization becomes computationally demanding but remains feasible for benchmark comparison.
Figure 7: Expectation values of the averaged magnetization components $\langle\sigma_{x}\rangle$, $\langle\sigma_{y}\rangle$, and $\langle\sigma_{z}\rangle$ as functions of the normalized transverse field strength $g/\gamma$ for a 2D dissipative Ising model on a $2\times 3$ square lattice with periodic boundary conditions. The system parameters are $V=2\gamma$ with uniform dissipation rate $\gamma$ across all 6 sites. The red dots represent results from our transformer density operator ansatz. The transformer architecture successfully handles the increased complexity of 2D geometry, including enhanced coordination numbers and frustration effects compared to 1D systems.
As shown in Fig. 7, our transformer density operator ansatz demonstrates excellent accuracy in capturing the steady-state magnetization components $\langle\sigma_{x}\rangle$, $\langle\sigma_{y}\rangle$, and $\langle\sigma_{z}\rangle$ across the entire range of transverse field strengths. The exact curve is calculated via the iterative BiCGStab solver of NetKet and qutip. The 2D geometry introduces additional complexity through the increased coordination number (four nearest neighbors per site) and enhanced frustration effects, yet our transformer approach handles these challenges effectively. We also include the result for $2\times 2$ in the supplementary material D. For large 2D system, more parameters are required to reach a high precision.
The self-attention mechanism proves particularly advantageous in the 2D setting, as it naturally captures the long-range correlations that emerge from the interplay between the 2D geometry and dissipative dynamics.
Conclusion and discussion. — In this work, we introduce a transformer variational ansatz to efficiently encode and solve the steady states of dissipative quantum systems. By vectorizing the density operator into a many-body state, we embed the input configurations into feature vectors and then utilize the self-attention mechanisms to model long-range correlation, which is crucial in the open system. Numerical experiments on paradigmatic models, including one-dimensional dissipative transverse-field Ising and Heisenberg models, as well as a two-dimensional dissipative Ising model, show that our approach accurately reproduces the mixed steady states of various dissipative systems while maintaining a compact parameterization and substantial complexity advantage.
@@
Acknowledgments. — We acknowledge Di Luo, Filippo Vicentini, Yuan-Hang Zhang, and Chen Zhuo for beneficial communications. We thank Filippo Vicentini and Di Luo for sharing their codes with us. The numerical implementation of the variational transformer density operator ansatz was done using JAX. The variational quantum Monte Carlo and stochastic reconfiguration optimizers are available in NetKet. Z. J. and D. K. are supported by the National Research Foundation in Singapore and A*STAR under its CQT Bridging Grant, CQT- Return of PIs EOM YR1-10 Funding and CQT Young Researcher Career Development Grant.
Appendix A Steady state and the vectorization of density operator
In this section, we review the vectorization formalism of the density operator and its dynamics, a powerful framework for representing the steady states of dissipative open quantum systems. This approach has recently gained significant attention in studies of open-system quantum phases, weak and strong symmetries, tenfold classification, and other related topics (see, e.g., Kawabata et al. (2023); Bao et al. (2023); Sohal and Prem (2025); Yoshioka and Hamazaki (2019); Ma and Turzillo (2024); Nigro (2020, 2019)). Unlike the traditional density matrix representation, the vectorized form offers a more convenient computational framework for steady-state analysis. By expressing the density operator in this form, steady states can be obtained by solving for the ground state of a specially constructed operator.
@@ Appendix A Steady state and the vectorization of density operator
$\hat{\rho}=\sum_{\alpha,\beta}\rho_{\alpha,\beta}|\alpha\rangle\langle\beta|,$ (18)
@@
$|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}=\sum_{\alpha,\beta}\rho_{\alpha,\beta}|\alpha\rangle|\beta\rangle.$ (19)
@@
$|(|\psi\rangle\langle\phi|)\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}=|\psi\rangle|\phi^{*}\rangle,$ (20)
@@
$|A\otimes B\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}\neq|A\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}\otimes|B\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}.$⟩ . (21)
@@
$A\rho B\mapsto(A\otimes B^{T})|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}.$⟩ . (22)
@@
$|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}=\sum_{\alpha_{1},\dots,\alpha_{N};\beta_{1},\dots,\beta_{N}}\rho_{\alpha_{1},\dots,\alpha_{N};\beta_{1},\dots,\beta_{N}}|\alpha_{1},\dots,\alpha_{N}\rangle\otimes|\beta_{1},\dots,\beta_{N}\rangle.$ (23)
@@
$\frac{d\rho}{dt}=\mathcal{L}(\rho)={-i}[H,\rho]+\sum_{i>0}\gamma_{i}(L_{i}\rho L_{i}^{\dagger}-\frac{1}{2}\{L_{i}^{\dagger}L_{i},\rho\}),$ (24)
@@
$\hat{\mathcal{L}}=$ $-i(H\otimes\mathds{I}-\mathds{I}\otimes H^{T})$ (26)
@@
$\mathfrak{L}=\hat{\mathcal{L}}^{\dagger}\hat{\mathcal{L}},$ (28)
@@
provides the steady state, as shown in equation (27). We can then optimize the energy functional to find the ground state, akin to the approach in closed systems. The energy functional is given by
$E=\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\rho|\mathfrak{L}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}=\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\rho|\hat{\mathcal{L}}^{\dagger}\hat{\mathcal{L}}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}.$⟩ . (30)
@@
$\operatorname{Tr}(A^{\dagger}B)=\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}A|B\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}.$⟩ . (31)
@@
$\langle O\rangle=\operatorname{Tr}(O\rho)=\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}O|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}},$⟩ , (32)
@@
Appendix B Transformer Density Operator Architecture
@@
B.0.1 Overview of the Transformer Density Operator Architecture
@@ B.0.2 Feature Embedding
$\mathbf{X}^{(1)}\;=\;\mathrm{Conv}_{\text{circular}}\!\bigl{(}\mathbf{X}_{\mathrm{input}}\bigr{)}\;\in\;\mathbb{R}^{B\times L\times 1\times C_{1}},$ (33)
@@
$\mathbf{X}^{(2)}\;=\;\mathrm{Conv}_{\text{circular}}\!\bigl{(}\mathbf{X}^{(1)}\bigr{)}\;\in\;\mathbb{R}^{B\times L\times 1\times C_{2}}.$ (34)
@@
B.0.3 Self-Attention for Global Correlations
@@
$\mathbf{X}\;=\;\mathrm{squeeze}\!\bigl{(}\mathbf{X}^{(2)},\,\texttt{axis}=2\bigr{)}\;\in\;\mathbb{R}^{B\times L\times C_{2}}.$ (35)
@@
$\mathbf{Q}$ $=\mathbf{X}\,\mathbf{W}^{Q},\quad\mathbf{K}\;=\;\mathbf{X}\,\mathbf{W}^{K},\quad\mathbf{V}\;=\;\mathbf{X}\,\mathbf{W}^{V},$ (36)
@@
$\mathrm{head}_{i}\;=\;\mathrm{softmax}\!\Bigl{(}\frac{\mathbf{Q}_{i}\,\mathbf{K}_{i}^{\top}}{\sqrt{C_{2}/h}}\Bigr{)}\,\mathbf{V}_{i}$ (37)
@@
$\mathbf{X}^{\prime}\;=\;\mathrm{Concat}\big{(}\mathrm{head}_{1},\dots,\mathrm{head}_{h}\big{)}\,\mathbf{W}^{O}\;+\;\mathbf{X}.$ (38)
@@
B.0.4 Global Mean-Pooling and Final Output
@@
$\mathbf{X}^{(\mathrm{pool})}\;=\;\frac{1}{L}\sum_{j=1}^{L}\bigl{(}\mathbf{X}^{\prime}_{:\,,j\,,:\,}\bigr{)},\quad\mathbf{X}^{(\mathrm{pool})}\,\in\,\mathbb{R}^{B\times C_{2}}.$ (39)
@@
Appendix C Optimization and Evaluation of Transformer Density Operator Ansatz
@@
C.1 Efficient Sampling Strategy
@@
$|\rho_{\bm{\theta}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}=\sum_{\bm{\alpha},\bm{\beta}}\rho_{\bm{\theta}}(\bm{\alpha},\bm{\beta})|\bm{\alpha}\rangle\otimes|\bm{\beta}\rangle.$ (40)
@@
$P_{\bm{\theta}}(\bm{\alpha},\bm{\beta})\propto|\rho_{\bm{\theta}}(\bm{\alpha},\bm{\beta})|^{2}$ (41)
@@
    $\mathbf{A}_{\textbf{accept}}((\bm{\alpha},\bm{\beta})\to(\bm{\alpha}^{\prime},\bm{\beta}^{\prime}))$ (42)
    $=$ $\min\left(1,~{}\frac{P_{\bm{\theta}}(\bm{\alpha}^{\prime},\bm{\beta}^{\prime})g((\bm{\alpha},\bm{\beta})|(\bm{\alpha}^{\prime},\bm{\beta}^{\prime}))}{P_{\bm{\theta}}(\bm{\alpha},\bm{\beta})g((\bm{\alpha}^{\prime},\bm{\beta}^{\prime})|(\bm{\alpha},\bm{\beta}))}\right),$
@@
C.2 Mixed-State Observables